OpenStudy (anonymous):

Which is less than 0.35 three-tenths eleven twenty-fifths two-fifths nine-twentieths

ganeshie8 (ganeshie8):

\(0.35 = \frac{35}{100}\) \(\frac{3}{10} = \frac{3*10}{10*10} = \frac{30}{100}\) \(\frac{11}{25} = \frac{11*4}{25*4} = \frac{44}{100}\) \(\frac{2}{5} = \frac{2*20}{5*20} = \frac{40}{100}\) \(\frac{9}{20} = \frac{9*5}{20*5} = \frac{45}{100}\)

OpenStudy (anonymous):

great!!

ganeshie8 (ganeshie8):

:) which one is it

OpenStudy (anonymous):

is it the first ome?

ganeshie8 (ganeshie8):

thats right ! just convert to same denominator so it become easy... its pain, but it becomes easy to figure out

OpenStudy (anonymous):

Thanks alot!

ganeshie8 (ganeshie8):

np... yw :D